# A Survey on Memory-Efficient Large-Scale Model Training in AI for Science

### v1

@@ IV Tailored Memory-Efficient Training Techniques: A Case Study
OpenFold [48] is an open-source implementation of AlphaFold 2, which has a prediction quality that matches AlphaFold 2 and runs faster on most proteins with less memory usage. This improves the ease and performance of training new models and performing large-scale predictions. OpenFold has made various optimizations to AlphaFold 2’s training schedule. OpenFold was trained on a cluster of 44 NVIDIA A100 GPUs, each with 40GB of memory. It enables FP16 mixed precision training mode to enhance training speed while minimizing memory usage. To reproduce the effective batch size utilized during AlphaFold 2 training as closely as possible, OpenFold implements 3-way gradient accumulation. In terms of distributed training, Deepspeed [85] and ZeRO stage 2 [75] are employed to distribute the optimizer states across each GPU, thus reducing model redundancy. Additionally, they refactored the model by replacing element-wise operations with in-place equivalents wherever possible to minimize unnecessary memory allocation. Advanced implementation of the self-attention [118], [119] was also adopted to lower memory usage and speed up attention computation. Furthermore, they optimize CUDA kernels for certain model modules, decreasing memory overhead.
@@
Unlike many existing LLMs, AlphaFold 2’s architecture derives its main memory overhead from intermediate activations, requiring more than 20 GB memory in each attention module, which cannot be well distributed using common methods like ZeRO DP, PP, or TP. Therefore, FastFold introduces a model parallel approach that focuses on reducing activations to optimize AlphaFold 2 training. Dynamic Axial Parallelism (DAP) is a technique tailored for the EvoFormer modules in AlphaFold 2. It retains complete model parameters on each device while partitioning inputs and intermediate activations across devices. The MSA representation and Pair representation within the EvoFormer module contain two distinct sequence dimensions, with computations performed along only one of these dimensions at a time. Consequently, DAP considers partitioning the other sequence dimension not in use. Due to the alternating calculations across different sequence dimensions, all-to-all communication is necessary to ensure that each device possesses the required data for computation. This design allows for the distribution of activations among devices, significantly reducing the training memory overhead of AlphaFold 2 and enhancing both efficiency and scalability. When the model is trained with DAP on a considerable number of devices, throughput can be improved by disabling gradient checkpointing, as DAP has saved enough memory. Compared with the original AlphaFold 2, FastFold has a training efficiency of 3.91×, and compared with OpenFold, it has improved by 2.98×.
@@
V Future Trends and Challenges
@@
The number of parameters and the data processing capabilities of large-scale models play a crucial role in their ability to address complex scientific problems. However, the memory requirements for training these models often surpass current hardware limitations, creating a significant bottleneck in scaling. Optimizing memory usage during training not only enhances efficiency but also provides possibilities for exploring even larger models, which may further improve performance across various tasks. This optimization is vital for scientific models, where tasks like protein structure prediction, molecular generation, and climate prediction demand models can handle high-dimensional and long-sequence data, while scaling up models can significantly enhance the accuracy and effectiveness of these applications. Moreover, employing effective memory optimization strategies allows practitioners to train models of the required size using fewer GPUs, resulting in substantial savings in cost and energy. It lowers the barriers to training large language models, promoting greater participation from researchers in scientific fields.
@@
For the training of transformer-based scientific large models, there exists diverse memory optimization methods and effective training libraries, such as DeepSpeed, Megatron-LM, Colossal-AI, FairScale, Hugging Face Accelerate, etc., all of which support distributed training, mixed precision training, gradient checkpointing, and other advanced techniques. However, our investigation indicates that many studies employing LLMs for scientific research have not fully leveraged these memory-efficient training approaches. Thus, there remains considerable room for optimization in large-scale model training in scientific tasks. This survey aims to systematically summarize the memory optimization techniques utilized in LLM training, hoping to promote their application in scientific domains and assist researchers in fully harnessing the capabilities of LLMs to accelerate scientific discovery and innovation.
@@
However, when it comes to large-scale models based on transformer variants, the memory optimization methods described in the paper, which are suitable for training standard Transformers, may not be fully applicable. Modules such as AlphaFold 2’s Evoformer, RosttaFold’s SE(3)-Transformer, and AlphaFold 3’s Pairformer depict notable differences in both memory usage and computational complexity compared to the vanilla Transformer architecture. Optimizing these models often requires domain expertise to develop tailored memory optimization strategies that align with the specific structure of the model. This process necessitates not only a profound understanding of the characteristics of corresponding scientific tasks but also the co-optimization of software and hardware to achieve optimal performance. Designing effective memory optimization methods for these models remains an urgent challenge that needs to be solved.
@@
As model size and task complexity continue to grow, memory-efficient training methods have become a key factor for advancing large-scale models for science. It is crucial to integrate memory optimization techniques with advanced training frameworks, develop customized optimization strategies, and further lower the resource requirements for model training. This represents a significant research direction. Furthermore, future research could focus on creating more versatile guidelines for designing optimization strategies and promoting the widespread use of high-performance computing resources in scientific fields through collaborative software and hardware design.
VI Conclusion
@@
In scientific fields, the performance and complexity of LLMs have continuously evolved as their applications expand into fields including biology, medicine, chemistry, and meteorology. However, training large-scale models often incurs significant memory demands and computational overhead, posing tough challenges to existing hardware resources. Memory-efficient training methods not only enhance training efficiency but also offer support for exploring the potential of larger-scale models by optimizing memory utilization, minimizing redundant computations, and improving resource allocation efficiency. This survey provides a systematic summary of the memory optimization techniques that have gained popularity in recent years for large-scale model training. Additionally, we use AlphaFold 2 as an illustrative example to introduce customized memory-efficient training approaches tailored to specific task requirements, serving as a beneficial reference for researchers interested in applications of large-scale models for scientific tasks.
@@
While existing memory optimization techniques and training frameworks for transformers have made significant strides in the field of NLP, their application in scientific fields has yet to achieve widespread adoption. Scientific models often encounter specific task requirements, including long-sequence modeling, high-precision numerical calculations, and multimodal data fusion. Therefore, it is essential to choose appropriate methods for memory optimization. Furthermore, future research should aim to develop more flexible and efficient optimization schemes tailored to these unique requirements. It is particularly crucial for large-scale models based on transformer variants, where memory optimization strategies must align with task characteristics and the model’s architecture. This will require researchers to make trade-offs between task demands, model design, and hardware implementation.

### v2

@@ 4 Tailored Memory-Efficient Training Techniques: A Case Study
OpenFold [43] is an open-source implementation of AlphaFold 2, which has a prediction quality that matches AlphaFold 2 and runs faster on most proteins with less memory usage. This improves the ease and performance of training new models and performing large-scale predictions. OpenFold has made various optimizations to AlphaFold 2’s training schedule. OpenFold was trained on a cluster of 44 NVIDIA A100 GPUs, each with 40GB of memory. It enables FP16 mixed precision training mode to enhance training speed while minimizing memory usage. To reproduce the effective batch size utilized during AlphaFold 2 training as closely as possible, OpenFold implements 3-way gradient accumulation. In terms of distributed training, Deepspeed [122] and ZeRO stage 2 [95] are employed to distribute the optimizer states across each GPU, thus reducing model redundancy. Additionally, they refactored the model by replacing element-wise operations with in-place equivalents wherever possible to minimize unnecessary memory allocation. Advanced implementation of the self-attention [130, 112] was also adopted to lower memory usage and speed up attention computation. Furthermore, they optimize CUDA kernels for certain model modules, decreasing memory overhead.
@@
Unlike many existing LLMs, AlphaFold 2’s architecture derives its main memory overhead from intermediate activations, requiring more than 20 GB memory in each attention module [44], which cannot be well distributed using common methods like ZeRO DP, PP, or TP. Therefore, FastFold introduces a model parallel approach that focuses on reducing activations to optimize AlphaFold 2 training. Dynamic Axial Parallelism (DAP) is a technique tailored for the EvoFormer modules in AlphaFold 2. It retains complete model parameters on each device while partitioning inputs and intermediate activations across devices. The MSA representation and Pair representation within the EvoFormer module contain two distinct sequence dimensions, with computations performed along only one of these dimensions at a time. Consequently, DAP considers partitioning the other sequence dimension not in use. Due to the alternating calculations across different sequence dimensions, all-to-all communication is necessary to ensure that each device possesses the required data for computation. This design allows for the distribution of activations among devices, significantly reducing the training memory overhead of AlphaFold 2 and enhancing both efficiency and scalability. When the model is trained with DAP on a considerable number of devices, throughput can be improved by disabling gradient checkpointing, as DAP has saved enough memory. Compared with the original AlphaFold 2, FastFold has reduced total training time from 11 days to 2.81 days, representing a reduction of 3.91×, and compared with OpenFold, it has improved by 2.98×.
@@
5 Future Trends and Challenges
@@
  • Scaling transformer models remains both a challenge and an opportunity. The number of parameters and the data processing capabilities of transformer models play a crucial role in their ability to address complex scientific problems. However, the memory demands for training these models often surpass current hardware limitations, creating a bottleneck in scaling. Optimizing memory usage during training not only enhances efficiency but also provides possibilities for exploring even larger models. This optimization is vital for scientific models, where tasks like protein structure prediction, molecular generation, and climate prediction demand models can handle high-dimensional and long-sequence data while scaling up models can enhance the accuracy of these applications. Moreover, employing effective memory optimization strategies allows practitioners to train models using fewer GPUs, resulting in substantial savings in cost and energy. It lowers the barriers to training large transformer models, promoting greater participation from researchers in scientific fields.
@@
  • Advanced memory optimization toolkits are available, but underutilized in science-focused LLM work. For the training of transformer-based scientific large models, there exists diverse memory optimization methods and training libraries, such as DeepSpeed, Megatron-LM, Colossal-AI, FairScale, Hugging Face Accelerate, etc., most of which support various optimization techniques introduced in this survey. However, our investigation indicates that many studies employing LLMs for scientific research have not fully leveraged these memory-efficient training approaches. Thus, there remains considerable room for optimization in LLM training in scientific tasks. This survey aims to systematically summarize the memory optimization techniques, hoping to promote their application in scientific domains and assist researchers in fully harnessing the capabilities of LLMs to accelerate scientific discovery and innovation.
@@
  • The optimization of transformer variants in science require customized optimization strategies. While many memory optimization techniques are effective for standard transformers, they may not generalize to modules like: Evoformer in AlphaFold 2, SE(3)-Transformer in RoseTTAFold, and Pairformer in AlphaFold 3, which deviate significantly in computational complexity and memory behavior. Optimizing these models often requires domain expertise to develop tailored strategies that align with the specific structure of the model. This process necessitates not only a profound understanding of the characteristics of corresponding scientific tasks but also the co-optimization of software and hardware to achieve optimal performance. Designing effective memory optimization methods for these models remains an urgent challenge that needs to be solved.
@@
  • Memory-efficient training is an essential direction for future research. As model size and data scale continue to grow, memory-efficient training methods have become a key factor for advancing transformer models for science. It is crucial to integrate memory optimization techniques with advanced training frameworks, develop customized optimization strategies, and further lower the resource requirements for model training. Furthermore, future research could focus on creating more versatile guidelines for designing optimization strategies and promoting the widespread use of high-performance computing resources in scientific fields through collaborative software and hardware design.
6 Conclusion
@@
In scientific fields, the performance and complexity of large transformer models have continually evolved as their applications expand into fields including biology, medicine, chemistry, and meteorology. However, their training often incurs significant memory demands and computational overhead, posing tough challenges to existing hardware resources. Memory-efficient training methods not only enhance training efficiency but also provide support for exploring the potential of larger-scale models in AI for science. This survey provides a systematic summary of the memory optimization techniques that have gained popularity in recent years for training transformers. Additionally, we use AlphaFold 2 as an illustrative example to introduce customized memory-efficient training approaches tailored to specific task requirements, serving as a beneficial reference for researchers interested in applications of transformer-based models for scientific tasks.
@@
While existing memory optimization techniques and training frameworks for transformers have made significant strides in NLP, their application in scientific fields has yet to achieve widespread adoption. Scientific models often encounter specific task requirements, including long-sequence input and multimodal data, which can result in high activation memory overhead, making it crucial to choose appropriate methods for memory optimization.